AliT06:

Ellen wants to purchase a book that is regularly priced at $18. The book is discounted 15%. She also needs to pay a 6% sales tax on the discounted price. Let me know is you are online I need tutoring

AliT06:

Answer when possible..

Nnesha:

There are two ways to solve this question: regular price is `$18` and the book is on sale `15%` off if you want to check how much money you will save then you can multiply \[\large\rm regular ~price *\frac{ percent ~ off}{100}\] convert 15% to decimal by dividing 15/100 \[18* 0.15= $2.70\] you will save $2.70 `18-2.70=$15.30` and then you have to pay 6% tax of 15.30 same formula again \[\large\rm price * \frac{percent }{100}\] `15.30 *0.06=0.918=$0.92` (6/100=0.06) 0.918 =0.92 rounded to two decimal places you have to pay 0.92 cents tax for %15.30 book so add 15.30+0.92= `$16.22` your total will be 16 dollar 22 cents. ~~~~~~~~~ 2nd way) if you wanna know how much money you will be paying: 1st) subtract the percent off from 100 (100-15=85%) 2nd) divide the difference by 100 85/100= 0.85 3rd) multiply the regular price by 0.85 `0.85*18=$15.30` 4th) when you have to add some percent then you have to add 100+6=106 5) divide 106 by 100 (106/100=1.06) 6)multiply it by the discounted price (15.30*1.06=16.218) 7) round it to two decimal places 16.218=` $16.22 ` same answer:=))

Nnesha:

let me know if that make no sense.